Directrix (directrix):

Given sin(x°) = .8, why is the Wolf giving 57.296 approx as the right triangle value for x when...

Directrix (directrix):

... the answer should be 53.130 degrees approx ? The Wolf's calculation here: http://www.wolframalpha.com/input/?i=sin%28+x+degrees%29+%3D+.8

OpenStudy (alexandervonhumboldt2):

impossible

OpenStudy (mrnood):

not quite sure what your issue is: If you hover over the point on the graph in your link where sin x = .8 it gives a value of 53.13 deg (note - the angle oyu give is 'almost' the complement of 53.13)

Directrix (directrix):

I understand and appreciate the postings of how I should have entered the query. What I don't follow is what question the answer of the 57ish degrees ( Reminds me of 1 radian.) In short, how is the "Wolf" interpreting my question as originally entered? http://www.wolframalpha.com/input/?i=sin%28+x+degrees%29+%3D+.8

OpenStudy (mathmath333):

its not only 57.26 its approx 57.26(6.28n+2.21) and 57.26(6.28n+0.92) for if u substitute n=1 \(\large\tt \begin{align} \color{black}{57.26(6.2823\cdot n+0.927)\hspace{.33em}\\~\\ =57.26(6.2823+0.927)\hspace{.33em}\\~\\ =412 =412.272-360 =52.272 \approx 53 }\end{align}\)

OpenStudy (asnaseer):

If you click on the "approximate forms" in the answer it displays two possible sets of solutions:\[x\approx57.296(6.2832n+2.2143)\]and\[x\approx57.296(6.2832n+0.9273)\]If you look at the second set of solutions and use \(n=0\) you will get \(x\approx5.13^{\circ}\)

OpenStudy (asnaseer):

Sorry - typo - I meant you will get \(x\approx53.13^{\circ}\)